関東女子タッチフットボール連盟役員

平成７〜８年度

理事長：冨田勝理事：阿部丈夫、大部善之、黒田恭司、田中憲一、冨田有子、吉川理裕運営委員：守屋聡子 (10/15 責任者)、内藤文子 (10/22 責任者)、山本貴子 (10/29 責任者)、坂本智子 (11/3 責任者)、広瀬和香子 (11/12 責任者)、内田好美、鍋島結子、太田雅子、渡辺知美、村上綾、中村浩子、牧野千代、千田若菜、原田理子、加藤正、中根弓枝、吉田有華、岩澤弥生、押川奈津美出版担当：吉田恵美子○、矢野敬子、宮崎奈々子記録連絡担当：牧田友子○、井上舞、吉次恭子渉外担当：三木淳子○、池田昌子会計担当：中川優子用度担当：植村萌子○、伊藤美咲、寺島佐知子、佐々木英美審判部長：森強審判員：秋田俊男、石川英司、須藤広幸、富田明、中西誠、萩原幸治、大和祥恭、吉村教幸、村上正秀、渡辺一夫準審判員：四ツ井亜希○、吉田恵美子、矢野敬子、峯岸まり子、中村浩子○、村上綾、山本貴子、植村萌子○、高田抄代子、亀山美紀、井上舞○、牧田友子、吉次恭子、牧野千代 ○、竹内香織、千田若菜、吉田有華○、岩澤弥生、諏訪綾子、内藤文子○、太田雅子、渡辺知美、原田理子○、中川優子、広瀬和香子、野沢厚子、冨田有子、細井智美、内田好美○、鍋島結子、守屋聡子

2

○は主任。